OpenStudy (anonymous):

how to do this ? What is the value of f(7) for the function f(x) = 4x + 3?

OpenStudy (anonymous):

plug in 7 at each place where is x

OpenStudy (anonymous):

47+3

OpenStudy (anonymous):

?

OpenStudy (phi):

yes, except 4x means 4 times x so 4x+3 with x replaced by 7 means 4*x+3

OpenStudy (phi):

4*7+3

OpenStudy (anonymous):

31

OpenStudy (anonymous):

how do i do this one What is the value of f[g(3)] for the functions f(x) = x + 3 and g(x) = 2x + 1?

OpenStudy (phi):

first do g(3), what do you get ?

OpenStudy (anonymous):

idnt know how to set this one up

OpenStudy (phi):

g(x) = 2x + 1 find g(3) means replace x with 3 (and remember 2x means 2*x)

OpenStudy (anonymous):

3 + 3 and 2 x 3 +1 right

OpenStudy (phi):

what did you get for g(3) ?

OpenStudy (anonymous):

what do u mean

OpenStudy (phi):

g(x) = 2x + 1 find g(3)

OpenStudy (anonymous):

confused how do i do that

OpenStudy (anonymous):

i thought g was 3

OpenStudy (phi):

g(3) is short hand for " everywhere you see x in 2x+1, put in a 3 instead"

OpenStudy (anonymous):

so 2*3 + 1

OpenStudy (phi):

yes, but simplify that to get a number

OpenStudy (phi):

what is 2 times 3 ? what is 2 times 3 and then add 1 ?

OpenStudy (anonymous):

6 and add the one is 7

OpenStudy (phi):

so g(3)=7 now f( g(3) ) is the same as f(7) (we replace g(3) with 7) what is f(7) ? remember f(x) = x + 3

OpenStudy (anonymous):

so it would b 7 + 3?

OpenStudy (anonymous):

i got confused on that part srry

OpenStudy (phi):

yes that is right f(7) = 7+3 what is 7+3 ?

OpenStudy (anonymous):

10

OpenStudy (phi):

that is the answer f(g(3))= 10

OpenStudy (anonymous):

thnks man

OpenStudy (anonymous):

what about this one Given the domain values { -2, 0, 2 }, find the range values for the function y = 5x + 1.

OpenStudy (anonymous):

dnt get it ?

OpenStudy (phi):

you should make a new post for each question.

OpenStudy (anonymous):

ok